OpenStudy (anonymous):

I just want to make sure i have this problem right. It will be posted below.

OpenStudy (anonymous):

\[\sqrt{x-3}\ge7\] I got\[3\le x \le 52\] or is it just \[x \ge 52\]

OpenStudy (kva1992):

the second one

OpenStudy (kva1992):

because you added the 3 to the other side thus removing the 3 and leaving the x

OpenStudy (solomonzelman):

it is just x≥52

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sqrt{x-3}≥7 }\) square both sides, \(\normalsize\color{blue}{ x-3≥49 }\) \(\normalsize\color{blue}{ x≥52 }\)

OpenStudy (kva1992):

this would be the only time you would get the first answer you posted \[0 \le x-3 \le 49\]

OpenStudy (kva1992):

hope i was able to help

OpenStudy (solomonzelman):

Bree Bree, you just confused it a little with solving similar equations with absolute values, where the x solutions (or the domain) is limited to he smallest AND biggest value that it can equal. Not just limited one way, and going infinitely to the other, like in a case of a square root .